# Get Your Embedding Space in Order: Domain-Adaptive Regression for Forest Monitoring

### v1

@@ 5.4 Ablation studies
Table 4: Ablation study for GOL+MDR. Average $R^{2}$ score across two source domains and three target domains. We report the performance of a model with only the center loss and weighted kNN prediction as the baseline.
@@
We further test the sensitivity of different methods to the number of shots (Supplementary Fig. 5), i.e., the number of support samples per group. As often observed in few-shot setups, performance saturates after a certain number of shots [38, 44], here 20. Our proposed GOL+MDR outperforms the other methods on most of the setups, except for the easiest task of tree cover, where finetuning is needed to reach comparable performance as other baselines.
@@ 6 Conclusion
We hope that the DRIFT dataset will encourage researchers to propose new methods for domain adaptation in image-level regression. In future work, we plan on extending the comparison to semi-supervised approaches on the target domain, with a larger dataset covering more domains across other continents.
@@ 9 Biogeographical regions
Supplementary Fig. 2: Biogeographical region distribution reflects the diversity of biomes in the DRIFT dataset. Image credits: European Environment Agency. https://www.eea.europa.eu/data-and-maps/figures/biogeographical-regions-in-europe-2

### v2

@@ 5.4 Ablation Studies
Table 4: Ablation study for combining GOL with the proposed MDR. Average $R^{2}$ score across two source domains and three target domains. We report the performance of a model with only the center loss and weighted kNN prediction as the baseline.
@@
We test the sensitivity of different methods to the number of shots (Supplementary Fig. 5), i.e., the number of support samples per group.
As often observed in few-shot setups, performance saturates after a certain number of shots [48, 58], here 20. Our proposed GOL+MDR outperforms the other methods on most of the setups, except for the easiest task of tree cover, where finetuning is needed to reach comparable performance as other baselines.
An important parameter in our approach is $k_{v}$, the number of support samples used for weighted sum prediction after MDR. We tested a set of values and observed a safe range around $k_{v}=2N$ (Supplementary Fig. 6).
@@ 6 Conclusion
Acknowledgements
PC acknowledges support from the European Space Agency Climate Change Initiative (ESA-CCI) Biomass project (ESA ESRIN/ 4000123662) and RECCAP2 project 1190 (ESA ESRIN/ 4000123002/ 18/I-NB) and the ANR BMBF French German AI4FOREST (ANR-22-FAI1-0002 ).
MB and SL were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement no. 947757 TOFDRY) and a DFF Sapere Aude grant (no. 9064–00049B).
DG ackowledge support by the European Union’s Eurostars programme through the C-Trees project, grant number E114613.
XT acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement no. 947757 TOFDRY).
Image Credits. Figures 1, 2 and Supplementary Figures 1, 3, and 4 contain data from the Danish Agency for Data Supply and Infrastructure (2018), the BD ORTHO© product from the French Institute for Geography (2018-2020), the Orthophoto and ALS products from the Slovakian authorities GKÚ Bratislava and ÚGKK SR (2021), the PNOA product from the Spanish Institute of Geography (scne.es, 2020), the SkySat sensor from Planet labs PBC, and the Slovenian Environment Agency (ARSO, 2021).
Supplementary Materials
7 DRIFT dataset details
Tree counts. For Denmark, the subset was built from the predictions of [33]. Due to the relatively high reported counting performance, we used the image-level counts directly as count annotations. For the rest, we corrected CHM-derived tree counts using linear regression with 3-fold validation. Correction factors ($y=ax+b$) and final $R^{2}$ scores are reported in Supplementary Table 1.
@@ 7 DRIFT dataset details
Label distribution. For the Denmark subset, we picked images to have a roughly uniform distribution of canopy height due to a limited variation of height observed in our initial random image selection. For the France and Slovakia subsets, images were randomly picked without any filtering or visual inspection. For the Spain and Slovenia subsets, we purposely chose forest-dominated regions prior to random picking to ensure efficient coverage of forest landscapes. The exact sampling locations are indicated in Supplementary Fig. 2. As such, the DRIFT dataset covers diverse label distributions representing different real-world scenarios.
Labeling uncertainties. The raw LiDAR point clouds have a horizontal (position) error of $\textsc{std}<20\text{cm}$ and a vertical (height) error of $\textsc{std}<10\text{cm}$, which are mitigated when extracting patch-level aggregates. We directly extracted tree cover fraction and non-zero mean height, without additional uncertainty.
Comparison with existing vegetation regression datasets. We compare the DRIFT dataset to existing regression datasets for vegetation monitoring in Supplementary Table 2. None of these datasets relies on sub-meter imagery over an area large enough to include significant domain shifts. Additionally, DRIFT includes three target variables with different levels of semantic complexity.
Supplementary Table 2: Vegetation regression datasets.
Size Target variable Resolution Domain shift
Potapov et al. 2019 [50] 56k Cover, Height 30m No
Kattenborn et al. 2020 [24] 4k Cover 5cm No
Hudak et al. 2020 [19] 3.8k Biomass 30m No
Feng et al. 2022 [15] 24k Cover, Age 30m Yes
Robin et al. 2022 [52] 40k Greenness 30m Yes
Lang et al. 2022 [28] 68k Height 10m Yes
Comparison with existing image-level regression datasets. Image-level regression examples can be found in facial age estimation [32], aesthetical quality estimation [55], facial pose estimation [13]. There has been limited interest in cross-domain regression, however. We summarize popular benchmark datasets on image-level regression in Supplementary Table 3.
Supplementary Table 3: Image-level regression datasets in broader computer vision.
Size Target variable Domain shift
Zhang et al. 2017 [66] 20k Facial age No
Moschoglou et al. 2017 [44] 16k Facial age No
Huang et al. 2016 [65] 41k Facial age No
Ricanek et al. 2006 [51] 55k Facial age No
Lin et al. 2021 [40] 287k Facial age No
Kuprashevich et al. 2023 [27] 67k Facial age No
Hosu et al. 2024 [18] 6k Image quality No
Murray et al. 2012 [46] 255k Aesthetic score No
Data licenses
  1. 1.
    Denmark
    20cm ground resolution RGB aerial imagery from the Danish Agency for Data Supply and Infrastructure (SDFI), 2018. CHMs are from the Danmarks Højdemodel product by SDFI, CC BY 4.0 license.
  2. 2.
    France
    20cm ground resolution RGB aerial imagery from the French Institute of Geography (IGN), BD ORTHO© product, 2018-2020. CHMs derived from the LiDAR HD product. All Data is under the 2.0 Open License.
  3. 3.
    Slovakia
    20cm ground resolution RGB aerial imagery from GKÚ Bratislava, NLC, 2021, downloaded from the GeoPortal. CHMs derived from the national ALS campaign, credits to ÚGKK SR, CC BY 4.0 license.
  4. 4.
    Spain
    25cm ground resolution RGB aerial imagery and CHMs derived from the Spanish Institute of Geography (IGN), PNOA 2020, scne.es. Data is under the CC-BY 4.0 License.
  5. 5.
    Slovenia
    50cm ground resolution RGB satellite imagery bought from Planet, SkySat sensor, 2021. Any usage must be solely for noncommercial education or scientific research purposes, and publication in academic or scientific research journals. All such publications must include an attribution that clearly and conspicuously identifies Planet Labs PBC. CHMs derived from the national ALS campaign, open license with attribution: Ministry of the Environment and Spatial Planning, Slovenian Environment Agency (ARSO).
@@ 9 Biogeographical Regions
Supplementary Fig. 2: Biogeographical region distribution reflects the diversity of biomes in the DRIFT dataset. The red polygons illustrate the exact locations from which image patches were extracted to build the DRIFT dataset. Image credits: European Environment Agency. https://www.eea.europa.eu/data-and-maps/figures/biogeographical-regions-in-europe-2